# ShaRP: A Novel Feature Importance Framework for Ranking

### v3

@@ Abstract.
Algorithmic decisions in critical domains such as hiring, college admissions, and lending are often based on rankings. Because of the impact these decisions have on individuals, organizations, and population groups, there is a need to understand them: to help individuals improve their position in a ranking, design better ranking procedures, and check whether a procedure is legally compliant. In this paper, we present ShaRP —Shapley for Rankings and Preferences—a framework that explains the contributions of features to different aspects of a ranked outcome and is based on Shapley values. Using ShaRP , we show that even when the scoring function used by an algorithmic ranker is known and linear, the feature weights do not correspond to their Shapley value contribution. The contributions instead depend on the feature distributions and the subtle local interactions between the scoring features.
ShaRP   builds on the Quantitative Input Influence framework to compute the contributions of features for multiple —ranking specific— Quantities of Interest, including score, rank, pair-wise preference, and top-$k$. We show the results of an extensive experimental validation of ShaRP  using real and synthetic datasets. We demonstrate that feature importance can be computed efficiently, and that ShaRP  compares favorably to several prior local feature importance methods, in terms of both generality and quality of explanations. Among our results, we highlight a case study on the CS Rankings dataset. Contrary to expectation, we find that a strong track record in Systems research is much more important than AI research for placing a CS department among the top-10%.
ShaRP  is available as an open-source library at https://github.com/DataResponsibly/ShaRP and is already used in teaching.
@@
ccs: Information systems Data management systemsccs: Social and professional topics Socio-technical systemsccs: Human-centered computing
@@ 1. Introduction
There are two types of rankers: score-based and learned. In score-based ranking, a given set of candidates is sorted on a score, which is typically computed using a simple formula, such as a weighted sum of attribute values (Zehlike et al., 2022a). In supervised learning-to-rank, a preference-enriched set of candidates is used to train a model that predicts rankings of unseen candidates (Li, 2014). In this paper, we focus on score-based rankers, which are often preferred in critical domains based on the premise that they are easier to design, understand, and justify than learning-to-rank models (Berger et al., 2019). Score-based rankers are a prominent example of the so-called “interpretable models” (Rudin, 2019): the scoring function, such as $Y_{1}=0.4\times gpa+0.4\times sat+0.2\times essay$ in a college admissions scenario, is based on a (normative) a priori understanding of what makes for a good candidate. And yet, despite being syntactically “interpretable,” a score-based ranker may not be “explainable,” in that, the designer of the ranker or the decision-maker who uses it may be unable to accurately predict or understand its output ((Miller, 2019; Molnar, 2020)). We illustrate this with an example.
@@
This example shows that the weight of a scoring feature may not be indicative of its impact on the ranked outcome. Intuitively, this happens because rankings are, by definition, relative, while feature values, and scores computed from them, are absolute. That is, knowing the score of an item says little about the position of that item in the ranking relative to other items.
There has been substantial work on quantifying feature importance for classification ((Covert et al., 2020; Datta et al., 2016; Guidotti et al., 2019; Lundberg and Lee, 2017; Mohler et al., 2020; Ribeiro et al., 2016b; Strumbelj and Kononenko, 2010)), and one may be tempted to think that these methods can be reused to explain ranking. Unfortunately, this is not the case, because in classification and regression, predictions are being made independently for each item. For example, in binary classification, if all points happen to fall on the positive side of the decision boundary, then all will receive the positive outcome. In contrast, in selection and ranking, an outcome for some item v is not independent of the outcomes for items $\mathcal{D}\setminus\{\textbf{v}\}$. That is, only one item can occupy a specific rank position, and exactly $k$ items can appear at the top-$k$.
In summary, ranking is intrinsically different from classification or regression, and it has been treated differently from classification and regression in predictive contexts, as evidenced by the rich literature on learning-to-rank (Li, 2014). Fairness in ranking has also been treated differently from fairness in classification (Zehlike et al., 2022a, b). Similarly, interpretability methods have to be customized for ranking, as we do in this work. The key insight then is that the impact of the scoring features on the ranking must be measured in a way that captures the lack of independence between per-item outcomes.
1.1. Summary of Our Approach
In this paper, we focus on local feature-based explanations of ranked outcomes. By “local,” we mean that we will explain the outcome for a particular item or a pair of items, although we may sometimes aggregate explanations across items to provide a summary. By “ranked outcome” we meant an item’s score, rank, its presence or absence among the top-$k$ (for some value of $k$), and the relative order among a pair of items. For consistency with prior literature, we will refer to different kinds of outcomes as “quantities of interest”, or QoIs for short (Datta et al., 2016). Finally, by “feature-based” we mean that we will quantify the impact of a feature’s value on a QoI.
We build on the conceptual Quantitative Input Influence (QII) framework by Datta et al. (2016) to quantify feature importance in predictive classification based on Shapley values ((Shapley et al., 1953)), a canonical method of dividing revenue in cooperative games. QII explains the influence of an item’s features (players) on the classification outcome for that item (cooperative game). We use QII as the starting point because this framework flexibly and naturally accommodates new quantities of interest. We augment QII with several Quantities of Interest (QoI) that are appropriate for ranking, including score, rank, pair-wise preference, and top-$k$.
We now preview some of the local feature-based explanations of ranked outcomes generated with ShaRP  over CS Rankings, a real dataset that ranks 189 Computer Science departments in the US based on a normalized publication count of the faculty across 4 research areas: AI, Systems, Theory, and Interdisciplinary (Berger, 2023) (see Section 7.1 for a detailed description of the dataset and the ranker). We will definite the QoIs that correspond to these ranked outcomes formally in Section 4 and will explain how ShaRP  computes them in Section 5. Here, our goal is to highlight some interesting insights about the data that these explanations can surface and, further, to show that these insights can be very different depending on what kind of ranked outcome is being explained.
@@ 1.1. Summary of Our Approach
(a) Score QoI: Interdisciplinar is the most important feature, followed by AI and Systems, while Theory negatively impacts the score.
@@
Example 0.
Figure 2 presents a feature-based explanation for Texas A&M in CS Rankings, where it has a score of 3.941 and is ranked in position 34. Figures 2a and 2b show feature importance for the score and rank QoIs, respectively. Intuitively, the score QoI quantifies the impact of the features on the score, as compared to the expected (average) score. The average score in CS Ranking is 2.722, shown as $E[f(X)]=2.722$ in Figure 2a. Features shown in red in that figure have a positive impact: they increase the score of Texas A&M by the increment that is shown inside the arrow (e.g., +0.49 for “Inter.”), as compared to the average score. Features shown in blue have a negative impact; they decrease the score compared to the average (e.g.,-0.11 for “Theory”). The sum of feature impact values corresponds to the difference between the average score of 2.722 and Texas A&M’s score of 3.941, up to rounding.
Similarly, rank QoI quantifies the impact of the features on the rank, as compared to the expected rank (94.5 out of 189 in CS Ranking), but with an important notational difference: a higher score but a lower rank is desirable. For this reason, feature importance for “Systems” — the most impactful feature when it comes to improving the rank of Texas A&M — has a negative sign, but it is still shown in red, for consistency. Finally, note that the absolute value of feature importance is higher for rank than for score (e.g.,-20.78 for “Systems”) because it quantifies the actual displacement in the ranking.
According to Figure 2a, the feature “Interdisciplinary” is the most (positively) impactful when it comes to improving the score of Texas A&M, followed by “AI” and then by “Systems”, while “Theory” contributes negatively. In contrast, all features contribute to improving the rank of Texas A&M, with “Systems” being the most impactful.
Example 2 demonstrates that feature-based explanations are different for different outcomes. If the goal is to improve the score, then Texas A&M should invest in interdisciplinary and AI research (features with highest scores in Figure 2a). However, if the goal is to directly improve the rank, then it’s better to prioritize systems research. Why are the score and rank QoI explanations different? The reason is that a large change in score may not result in a comparable change in rank (and vice versa). On a related note, average score is influenced by outliers, while rank is not. In fact, over 70% of the departments in CS Rankings have scores below the mean, and, as a result, when using score QoI, many or all of their features appear to be contributing negatively. (Figure 2b shows the score vs. rank distribution for this dataset.)
Example 0.

### v4

# ShaRP: Explaining Rankings and Preferences with Shapley Values

@@ Abstract.
Algorithmic decisions in critical domains such as hiring, college admissions, and lending are often based on rankings. Given the impact of these decisions on individuals, organizations, and population groups, it is essential to understand them—to help individuals improve their ranking position, design better ranking procedures, and ensure legal compliance. In this paper, we argue that explainability methods for classification and regression, such as SHAP, are insufficient for ranking tasks, and present ShaRP—Shapley Values for Rankings and Preferences—a framework that explains the contributions of features to various aspects of a ranked outcome.
ShaRP computes feature contributions for various ranking-specific profit functions, such as rank and top-$k$, and also includes a novel Shapley value-based method for explaining pairwise preference outcomes. We provide a flexible implementation of ShaRP, capable of efficiently and comprehensively explaining ranked and pairwise outcomes over tabular data, in score-based ranking and learning-to-rank tasks. Finally, to evaluate ShaRP and compare it with other explainability methods, we define ranking-specific explanation metrics and conduct an extensive experimental analysis, demonstrating the framework’s flexibility and efficiency.
@@ 1. Introduction
There are two types of rankers: score-based and learned. In score-based ranking, a given set of candidates is sorted on a score, which is typically computed using a simple formula, such as a weighted sum of attribute values (Zehlike et al., 2022a). In supervised learning-to-rank (LtR), a preference-enriched set of candidates is used to train a model that predicts rankings of unseen candidates (Li, 2014). Score-based rankers are a prominent example of the so-called “interpretable models” (Rudin, 2019): the scoring function, such as $Y_{1}=0.4\times gpa+0.4\times sat+0.2\times essay$ in a college admissions scenario, is based on a (normative) a priori understanding of what makes for a good candidate. And yet, despite being syntactically “interpretable,” a score-based ranker may not be “explainable,” in that, the designer of the ranker or the decision-maker who uses it may be unable to accurately predict or understand its output ((Miller, 2019; Molnar, 2020)). We illustrate this with an example.
@@
This example illustrates that a scoring feature’s weight may not reflect its impact on the ranked outcome. This occurs because rankings are inherently relative, while feature values and computed scores are absolute—knowing an item’s score reveals little about its position relative to others. The lack of independence between per-item outcomes renders existing feature importance methods for classification and regression (Covert et al., 2020; Datta et al., 2016; Guidotti et al., 2019; Lundberg and Lee, 2017; Mohler et al., 2020; Ribeiro et al., 2016b; Strumbelj and Kononenko, 2010) insufficient for ranking tasks. These methods assess how a feature influences an item’s score, but a feature may significantly affect an item’s score while leaving its rank unchanged. Consider the following example.
Example 0.
Consider Figure 1 and suppose that Dia’s essay score increases from 4 to 5, thus increasing the scores computed with both $f$ (4.4 to 4.6) and $g$ (4 to 5). However, Dia’s rank remains unchanged.
Changes in score do not necessarily lead to changes in rank because, in selection and ranking, an item’s outcome v depends on the outcomes of other items in $\mathcal{D}\setminus\{\textbf{v}\}$. For example, only one item can occupy a given rank, and exactly $k$ items can appear at the top-$k$. Thus, any explainability method that measures score changes can only partially explain rank changes. This highlights that interpretability for ranking tasks requires measuring the features’ impact on quantities beyond the score, such as rank or top-$k$ presence. We preview these results for CS Rankings in Figure 2, where feature importance for score in 2a and rank in 2b yield markedly different explanations. We discuss these findings in detail in Section 4.
@@
(a) Score QoI: Interdisciplinary is the most important feature, followed by AI and Systems, while Theory negatively impacts the score.
@@
In summary,
ranking fundamentally differs from classification and regression, as reflected in the extensive literature on learning-to-rank (Li, 2014). Fairness in ranking has also been addressed separately from fairness in classification (Zehlike et al., 2022a, b). Likewise, interpretability methods must be tailored for ranking, as we do in this work. The key insight is that the impact of scoring features on ranking must be measured in a way that accounts for the lack of independence between per-item outcomes. This work formalizes and builds upon this intuition, making four contributions.
Our first contribution is the formalization of several natural profit functions for computing Shapley values in ranking, capturing feature contributions to an item’s score, rank, or top-$k$ presence
We build on the Quantitative Input Influence (QII) framework by Datta et al. (2016), which quantifies feature importance in classification using Shapley values (Shapley et al., 1953), a standard method for revenue division in cooperative games. In QII, an item’s features (players) influence its classification outcome (profit function) within a cooperative game. We adopt QII as a flexible foundation for exploring different profit functions in Shapley value explanations. QII refers to profit functions as Quantities of Interest (QoIs), a term we use in this work.
Our second contribution is a new Shapley value method for explaining pairwise outcomes (preference pairs).
A class of Shapley value methods generates explanations by comparing an item to a “baseline” item considered “neutral” (Lundberg and Lee, 2017; Chen et al., 2023). Instead of using a fixed baseline item, we dynamically adjust it for each pair $u\prec v$, computing feature importance for $v$ using $u$ and vice versa. This approach yields feature contributions that naturally explain the difference between the compared items’ outcomes.
Our third contribution is a robust and extensible open-source library called ShaRP , the only available library that uses Shapley values for explaining ranked outcomes over tabular data.
Our implementation includes both exact and approximate QoI computation to achieve interactive performance while retaining explanation quality. As part of ShaRP , we are also making an implementation of the original QII framework available to the community.
Our fourth contribution is a methodology for evaluating ranking explainability methods for tabular data, supported by extensive experiments.
We define evaluation metrics and conduct qualitative analyses on real and synthetic datasets, showing that feature importance for all QoIs depends on data distribution properties and ranking position, in addition to the scoring function. Additionally, we demonstrate that the rank-aware QoIs we introduce provide unique and complementary insights beyond score-rank relationships. We also present a large-scale quantitative evaluation of ShaRP , highlighting the scalability of approximations and comparing it to other methods.
2. Related Work
Local feature-based explanations
Ribeiro et al. (2016b) introduced LIME, which explains complex classification models locally using interpretable models like sparse linear regressions. Lundberg and Lee (2017) proposed SHAP, which leverages Shapley values to explain various classification and regression models. Both methods are extensively implemented in software libraries and can be used to explain an item’s score (score QoI in our terminology).
Feature-based explanations for ranking
Yang et al. (2018) introduced a “nutritional label” for score-based rankers, featuring two visual widgets: “Recipe,” which displays scoring feature weights, and “Ingredients,” which highlights features (scoring or not) with the strongest Spearman’s rank correlation with the score. They observed that a feature’s weight in “Recipe” often does not correspond to its correlation with the score. This approach provides global feature-based explanations, identifying which features influence the ranked outcome for all items. In contrast, we focus on local explanations, highlighting feature importance for a specific item or pair of items.
Gale and Marian (2020) introduced “participation metrics” for score-based rankers, measuring how scoring features contribute to an item’s presence in the top-$k$. Their key metric, “weighted participation,” attributes an item’s inclusion in the top-$k$ to its feature values, scoring function weights, distance from maximum possible values, and whether a feature value exceeds the $k^{th}$ highest value. This quantity is aggregated across all top-$k$ items, producing a single participation value per feature. The main distinction between our method and Gale and Marian (2020) is that we provide local (per-item) explanations, whereas their approach offers group-wise explanations for the top-$k$. Our top-$k$ QoI is similar to weighted participation but focuses on local explanations, which can optionally be aggregated across the top-$k$.
Yuan and Dasgupta (2023) developed an interface for sensitivity analysis of ranked synthetic datasets with two normally distributed features and a linear scoring function. They quantify feature importance by subtracting a feature’s value from its mean, which, under their assumptions, approximates Shapley values (Lundberg and Lee, 2017). They also propose two alternative methods for quantifying feature contributions using the linear weights and the item’s score/rank. To compare with their approach, we re-implemented and significantly extended their method to support more than two features, arbitrary feature distributions, and flexible scoring functions.
Anahideh and Mohabbati-Kalejahi (2022) provide local feature-based explanations using multiple methods, including Shapley values computed with SHAP (Lundberg and Lee, 2017), and apply these methods locally to items ranked around the one being explained. They justify this approach by noting that in repeating competitions, such as school rankings, items tend to maintain similar ranks. Like Anahideh and Mohabbati-Kalejahi (2022), we find that feature importance varies significantly across rank strata. However, contrary to their assumptions, we show that slight changes in feature values can cause substantial rank shifts, extending far beyond the local neighborhood.
Moskovitch et al. (2023) developed DEXER to detect groups that may be under- or overrepresented in the top-$k$ compared to their proportion in the input. To analyze the causes of these disparities, they fit a linear regression model to ranks (treating them as scores) and compute feature importance using SHAP (Lundberg and Lee, 2017). In contrast, we fully adapt Shapley values for rank-specific outcomes in our ShaRP  framework. We compare the explanations generated by ShaRP  and DEXER in Section 8.2.
Similarly, Pastor et al. (2021) proposed a method that uses itemsets to automatically detect under- or overrepresented subgroups, where each subgroup is defined by a subset of the attributes. Like our approach, they define a ranking profit function, but they use it to quantify each attribute’s contribution to the itemset’s outcomes.
Finally, Hu et al. (2022) introduced PrefSHAP, a method for explaining pairwise preferences of dueling data modeled using the generalized preferential kernel (GPM) from their prior work (Chau et al., 2022). GPM is ideal for modeling inconsistent preferences that do not lead to a global ranking. Similar to our paper, they show that score-based Shapley values are insufficient to explain GPM preferences, and they define a new Shapley value profit function for this type of data. To compute Shapley value contributions, they apply the transformations for every coalition on both items at the same time and evaluate how the preference changes. Like Hu et al. (2022), we argue that ranking-specific quantities of interest (QoIs) are essential for explaining rankings and preferences, and introduce a broader set, including score, rank, and top-$k$, and a new pairwise explainability method. We do not compare experimentally with PrefSHAP because our focus is on preferences induced by score-based rankers or LtR, whereas PrefSHAP applies only to a specific class of kernel models.
Shapley-based explanations in Information Retrieval (IR)
Concurrently with our work, Heuss et al. (2024) and Chowdhury et al. (2024) published preprints adapting Shapley values for explaining ranked outcomes in IR. Both methods generate feature importance for the entire ranking rather than individual ranked items, therefore producing global Shapley-based explanations. To achieve this, they use SHAP (Lundberg and Lee, 2016), varying the same subset of features in each coalition for all items (documents). In contrast, our explanations are local (explaining outcomes for one item at a time), and so our work is not directly comparable to (Heuss et al., 2024; Chowdhury et al., 2024). Furthermore, our focus is on explaining rankings over tabular data, rather than over document collections. Other recent work in IR explored the use of LIME to explain ranked outcomes (Verma and Ganguly, 2019; Singh and Anand, 2018; Chowdhury et al., 2023), and introduced baseline document construction techniques for DeepSHAP (Lundberg and Lee, 2017) to improve explanation quality (Singh and Anand, 6 29).
In summary,
we share motivation with all these lines of work but take a leap, describing the first comprehensive Shapley-value-based framework for explaining ranking and pairwise preference outcomes.
3. Preliminaries and Notation
Ranking.
A ranking $r_{\mathcal{D}}$ is a permutation over the items in $\mathcal{D}_{\mathit{}}$. Letting $n=\left|\mathcal{D}_{\mathit{}}\right|$, we denote by $r_{\mathcal{D}}=\langle\textbf{v}_{1},\ldots,\textbf{v}_{n}\rangle$ a ranking that places item $\textbf{v}_{i}$ at rank $i$. We denote by $r_{\mathcal{D}}(i)$ the item at rank $i$, and by $r_{\mathcal{D}}^{-1}(\textbf{v})$ the rank of item v in $r_{\mathcal{D}}$. In score-based ranking, we are interested in rankings induced by some scoring function $f$. We denote these rankings $r_{\mathcal{D},f}$. For example, in Figure 1b, $r_{\mathcal{D},f}(1)=\texttt{Bob}$, $r_{\mathcal{D},f}^{-1}(\texttt{Leo})=7$. We assume that $r_{\mathcal{D},f}^{-1}(\textbf{v}_{1})<r_{\mathcal{D},f}^{-1}(\textbf{v}_{2})<\dots<r_{\mathcal{D},f}^{-1}(\textbf{v}_{n})$, where smaller rank means better position in the ranking.
[…]
Shapley values
We will use Shapley values to explain ranked outcomes using the set of features $\mathcal{A}$ as the players, and the outcome (or the quantity of interest, QoI) as the payoff function. In addition to the definition of players and the payoff function, Shapley values require the quantification of the payoff over a subset of the players. This, in turn, requires some way to estimate the payoff over a subset of the features. Consequently, for any Shapley value implementation, a method of feature removal or masking is required (Covert et al., 2021; Chen et al., 2023).
A common method (e.g., used in SHAP (Lundberg and Lee, 2017; Covert et al., 2021)), for a coalition $\mathcal{S}\subseteq\mathcal{A}$, is to marginalize out the features not in the coalition $\mathcal{A}\setminus\mathcal{S}$ and draw values from the marginal distributions of the subset of features in $\mathcal{S}$ jointly, often referred to as the “marginal” approach. Another alternative (e.g., used in QII (Datta et al., 2016)) is to draw values of each feature in $\mathcal{S}$ independently from its marginal distribution, often referred to as the “product of marginals” approach. Another approach is called “baseline” and instead of sampling the features in the coalition, they are replaced with the feature values of a specific fixed sample (Lundberg and Lee, 2017). Here, we choose the marginal approach for our implementation and take inspiration from the baseline approach for our pairwise method. In Section 7, we show how both can be implemented using one algorithm.
Let $\textbf{v}_{\mathcal{S}}$ denote a projection of v onto $\mathcal{S}$. In the example in Figure 1, $(\texttt{Bob},\texttt{4},\texttt{5},\texttt{2})_{\{name,gpa\}}=(\texttt{Bob},\texttt{4})$. We draw values from the marginal distributions of the subset of features in $\mathcal{S}$ as follows. Let U $=\langle\textbf{u}_{1},\dots,\textbf{u}_{m}\rangle$ denote a random vector of $m$ items sampled from $\mathcal{D}$. For a subset of features $\mathcal{S}\in\mathcal{A}$, let $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}=\langle\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}(\textbf{u}_{1})_{\mathcal{S}},\dots,\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}(\textbf{u}_{m})_{\mathcal{S}}\rangle$ denote a vector of items, in which each $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}(\textbf{u}_{i})_{\mathcal{S}}$ takes on the values of the features in $\mathcal{S}$ from $\textbf{u}_{i}$, and the values of the remaining features $\mathcal{A}\setminus\mathcal{S}$ from v. We calculate Shapley values using this set of features $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}$, note that if $m=|\mathcal{D}-1|$ we use the entire dataset $\mathcal{D}\setminus\textbf{v}$ to calculate the exact Shapley values.
Shapley values satisfy several natural axioms, including efficiency, symmetry, dummy, and additivity (Shapley et al., 1953), with additional useful properties, such as monotonicity, following from these axioms (Lundberg and Lee, 2017). Efficiency states that the sum of the contributions of all features for item v equals the difference between the outcome $f(\textbf{v})$ and the average outcome: $\sum_{i\in\mathcal{A}}\varphi_{i}(f,\textbf{v})=f(\textbf{v})-\mathbb{E}_{\textbf{X}}[f(\textbf{X})]$ (Covert et al., 2020; Molnar, 2020). Using this property, the explanation can be used to reconstruct the outcome. We will use the efficiency property to define the fidelity metric for comparing explanations (Section 6).
4. Quantities of Interest for Ranking
The first contribution of our work is that we define QoIs that are appropriate for ranked outcomes. In addition to the score, we introduce the rank and top-$k$ QoIs. We use the notation for the marginal feature removal approach in this section but note that the QoIs we introduce can be used with any feature removal approach.
Score QoI
The Shapley value function for the score QoI is:
This QoI captures the impact of an item’s features on its score. This is the QoI used by the popular feature-based explanation methods such as SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016a). To get the contribution of a set of features $\mathcal{A}\setminus\mathcal{S}$, we take the expected value of the score over $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}$ using the random variable $\textbf{U}_{\mathcal{S}}$ that takes values from the marginal distributions of the set of features in $\mathcal{S}$.
Rank QoI
The Shapley value payoff function for the rank QoI is:
where $\mathcal{D}^{\prime}$ is $\mathcal{D}\cup\{\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}\}\setminus\textbf{v}$. This QoI evaluates the impact of an item’s features on its rank. To get the contribution of a set of features $\mathcal{A}\setminus\mathcal{S}$, we take the expected value of the rank over $\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}$ using the random vector $\textbf{U}_{\mathcal{S}}$ using the random variable $\textbf{U}_{\mathcal{S}}$ that taakes values from the marginal distributions of the set of features in $\mathcal{S}$.
Top-$k$ QoI
(4) $QoI_{f,\textbf{v},\mathcal{D}}(S)=\mathop{\mathbb{E}}_{\textbf{U}_{\mathcal{S}}}[\mathbbm{1}_{r_{\mathcal{D}^{\prime},f}(1\ldots k)}(\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}})]$
where $\mathcal{D}^{\prime}$ is $\mathcal{D}\cup\{\textbf{v}_{\mathcal{A}\setminus\mathcal{S}}\textbf{U}_{\mathcal{S}}\}\setminus\textbf{v}$. The difference with rank QoI (Equation 3) is that here we compute the expectation over the indicator function that returns 1 if an item’s rank is at most $k$ and 0 otherwise. This QoI allows us to quantify how each feature contributed to getting the item into the top-$k$.
Shapley values for ranking
To compute Shapley values for the QoIs we defined, we need to apply Equation 1 on the QoIs. Following the QII notation, we define the iota function $\iota$ as the difference between the QoIs of $\mathcal{S}\cup i$ and $\mathcal{S}$, the coalition including feature $i$ and the coalition.
(5) $\iota_{f,\mathbf{v},\mathcal{D}}(i,\mathcal{S})=\alpha(QoI_{f,\mathbf{v},\mathcal{D}}(\mathcal{S}\cup i)-QoI_{f,\mathbf{v},\mathcal{D}}(\mathcal{S}))$
where, the QoI can be any of the QoIs defined in this section, and $\alpha\in\{-1,1\}$ is a multiplier that adjusts the order of QoI terms. This is necessary when consider QoIs such as rank, where smaller values are preferable. In this case, we set $\alpha=-1$ to adjust the $\iota$ function accordingly.
Using this notation, we can define Shapley values for ShaRP:
(6) $\phi_{i}(f,\mathbf{v},\mathcal{D})=\sum_{\mathcal{S}}\frac{|\mathcal{S}|!(n-|\mathcal{S}|-1)!}{n!}\iota_{f,\mathbf{v},\mathcal{D}}(i,\mathcal{S})$
Case Study: QoIs for CSRankings
We now review some of the local feature-based explanations of ranked outcomes generated with ShaRP  over CS Rankings, a real dataset that ranks 189 Computer Science departments in the US based on a normalized publication count of the faculty across 4 research areas: AI, Systems, Theory, and Interdisciplinary (Berger, 2023), see Appendix  A for a detailed description of the dataset and the ranker. Here, our goal is to highlight some interesting insights about the data that these explanations can surface and, further, to show that these insights can be very different depending on what kind of outcome is being explained.
In Figure 3, we show feature contributions to rank and score QoI for the CS Rankings dataset, aggregated over 10% strata of the ranking. According to rank QoI in Figure 3a Systems is the most important feature in all strata, followed closely by AI. Both features have the most positive contributions in the top strata and the most negative in the bottom strata. Further, we observe that feature contributions to score are less informative than to rank: both capture the same relative feature importance for the top 20%; however, feature contributions become small and very similar as more items are tied or nearly tied for their scores, making it nearly impossible to quantify feature importance across strata.